OpenStudy (anonymous):

find a fourt-dregree polynomial function passing through the point (-1,16) with roots of 0,1(multiplicity 2), and -3, i am studying for a quiz

OpenStudy (anonymous):

you know this problem?

OpenStudy (dumbcow):

f(x) = a*x(x+3)(x-1)^2 solve for a using given point 16 = a*(-1)(-1+3)(-1-1)^2 solve for a

OpenStudy (anonymous):

wait, a*x?

OpenStudy (dumbcow):

sorry, just means multiplying

OpenStudy (anonymous):

a*x?

OpenStudy (anonymous):

where you get a and x anyway?

OpenStudy (dumbcow):

a is like a scale factor, it determines how skinny or wide the graph is kind of like the slope

OpenStudy (anonymous):

so it should be x infrotn of (-1) etc?

OpenStudy (dumbcow):

no i was replacing x with -1 because the function passes through the point (-1,16) where the x_value is -1

OpenStudy (anonymous):

then -1(-1) etc, ? i got confused with ax , but i get it the you replace it.

OpenStudy (anonymous):

i understand what you have to do next, multy,

OpenStudy (dumbcow):

ah ok lets say you have a parabola y = (x-1)^2 it has one zero at x=1 right? but what about y = 5(x-1)^2 it also has one zero at x=1, however its a different function, its a narrower parabola so given the zeroes we can put the function into factored form but we dont know how the graph looks thats where the a comes from. the point given allows us to solve for a thus giving us the entire function

OpenStudy (dumbcow):

yes multiply and get a by itself

OpenStudy (anonymous):

yeah i see it, so is going to be , 16=-1(etc?

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

ok thx, going go to sleep, at what time you usually on dude?

OpenStudy (dumbcow):

depends, usually at night im on pacific time

OpenStudy (anonymous):

ok, so at this time?

OpenStudy (dumbcow):

yeah and definitely later too

OpenStudy (anonymous):

ok, bye

OpenStudy (anonymous):

hey, ok, you subtitude, then what?

OpenStudy (anonymous):

was stoke for awhile, while working on it

OpenStudy (dumbcow):

oh ok same one so you substituted the x,y values in from the point they gave you then you solve for a, get it by itself

OpenStudy (anonymous):

what i got right now 16=a(-1)(-1+3)(-1-1)

OpenStudy (dumbcow):

ok use PEMDAS do the parenthesis first (-1-1)^2 = (-2)^2 = 4

OpenStudy (anonymous):

ok now you got, 16=a(-1)(4)

OpenStudy (dumbcow):

what happened to the (-1+3) ??

OpenStudy (anonymous):

forogt..

OpenStudy (anonymous):

8

OpenStudy (dumbcow):

yep good 16 = a(-1)(8)

OpenStudy (anonymous):

16=-8a -2=a

OpenStudy (anonymous):

we found a, now we must replace a with -2

OpenStudy (dumbcow):

correct now use that and put it back in the equation

OpenStudy (anonymous):

to what equipon. 16=-2x(x+3)(x-1)^2?

OpenStudy (dumbcow):

the very first one now we put the x's and y's back in

OpenStudy (anonymous):

?

OpenStudy (dumbcow):

f(x) = ax(x+3)(x-1)^2

OpenStudy (dumbcow):

just replace a with -2

OpenStudy (anonymous):

-2a(1)(9) -18a

OpenStudy (anonymous):

how you find the fourt root?

OpenStudy (dumbcow):

woah not x, replace a with -2 thats what we just solved for

OpenStudy (dumbcow):

??? fourth root of what

OpenStudy (anonymous):

you looking for the fourt root

OpenStudy (dumbcow):

no its a fourth degree polynomial it just means our function has a x^4 term

OpenStudy (anonymous):

ok, how you start then?

OpenStudy (dumbcow):

i dont know what you mean you just did it answer is f(x) = -2x(x+3)(x-1)^2

OpenStudy (anonymous):

but the question is, find the fourt root

OpenStudy (anonymous):

oh NOT

OpenStudy (anonymous):

is find the fourt dree polynomial function

OpenStudy (anonymous):

i thought all along,to find the fourt root

OpenStudy (dumbcow):

there are 3 real roots 0,-3,1

OpenStudy (anonymous):

no really, their is four, 1 has multicplicity of 2

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

ok thank, yoiu being great help to me,good night

OpenStudy (dumbcow):

:)